Sylvia's monthly salary is $500.00 less than twice her daughter's salary. Together they earn $3,400 each month. How much is Sylvia's daughter's montly salary?
Answer by Maths68(1474)

Let
Daughter's monthly salary = x
Sylvia's monthly salary =2x-500
Given
Together they earn = $3,400
Daughter's monthly salary + Sylvia's monthly salary = 3400
x+(2x-500)=3400
x+2x-500=3400
3x-500=3400
3x=3400+500
3x=3900
Divide by 3 both sides of the equation
x=3900/3
x=1300
Daughter's monthly salary = $1300
Sylvia's monthly salary =2(1300)-500=2600-500=$2100